2 brothers have a combined age of 42 years. Steve is twice as old as Bob was when Steve was as old as Bob is now. How old are they?
Answer by TimothyLamb(4379)

x = steve's age now
y = bob's age now
---
x + y = 42
x = 2(y - (x - y))
---
put the system of linear equations into standard form
---
x + y = 42
x = 2(y - x + y)
---
x + y = 42
x = 2(2y - x)
---
x + y = 42
x = 4y - 2x
---
x + y = 42
3x - 4y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = steve's age now = 24
y = bob's age now = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php